Find the pH of the following: Show all work

  1. 0.020 moles  of  HCl  in 20.0 L of water
  2. 6 grams NaOH ( MW 40 g/mol)  in 15 L of water
  3. The titration  where Va= 5 mL   Vb= 20 L Ma= 0.40  Mb= ?
1
Expert's answer
2021-03-30T07:55:19-0400

"M= \frac{0.020}{20}=1.0×10^{-3}"


pH = -log (M)


pH= "-log(1.0\u00d710^{-3})= 3"



Question 2


"Moles = \frac{6}{40}=0.15 moles"


"M= \frac{0.15}{15}=0.01"


"pH=-log(0.01)=2"